# Alberta Wells Dataset: Pinpointing Oil and Gas Wells from Satellite Imagery

### v1

@@ A.4 Dataset Size & Distribution of Samples
A.5 PlanetScope Satellite Imagery
(a) A sample patch with Bbox annotations and the corresponding imagery in its different spectral bands.
(b) A sample patch with its segmentation labels (binary and multi-class) and bounding box annotations.
Figure 5: A Sample Patch from the Evaluation Set with 2 active wells.
For our experiments, we selected a 4-band (RGBN) satellite imagery product (ortho_analytic_4b_sr) from Planet Labs [16] as illustrated in Figure 5. This product uses Planet’s PSB.SD instrument, which features a telescope with a larger 47-megapixel sensor and is designed to be interoperable with Sentinel-2 imagery in several bands. The frequency of each band of image is described in Table 8. The instrument provides a frame size of 32.5 km x 19.6 km, an image capture capacity of 200 million km²/day, and an imagery bit depth of 12-bit, with a ground sample distance (nadir) ranging from 3.7 m to 4.2 m.
A.6 Meta Data Description
@@
A.7 Label Data Description
For our experiments, we create single-channel segmentation maps, which are binary maps used to locate instances of wells. We also generate multi-class segmentation maps, where each class denotes a well in an active, abandoned, or suspended state. Furthermore, we provide COCO format object detection labels for wells. In both segmentation and detection labels, we represent various states with class IDs as ’Active’: 1, ’Suspended’: 2, ’Abandoned’: 3. To maintain consistency, we standardize the diameter of a well site to 90 meters (typically ranging from 70 to 120 meters) when annotating, resulting in a 30-pixel diameter in the labels. Figure 5(b) illustrates image patches with their corresponding labels, Figure 5(a) illustrates various spectral bands present in an image alongside the original image with bounding box annotations for reference and an example of a bounding box label in COCO format is shown below.
@@ A.7 Label Data Description
Appendix B Challenges for ML Community
Appendix C Licenses
The AWD Dataset will be released under a Creative Commons Attribution-NonCommercial 4.0 International (CC BY-NC 4.0) License (https://creativecommons.org/licenses/by-nc/4.0/).The satellite imagery for this project was acquired through Planet Labs’ [16] Education & Research license, which allows the use of the data in publications and the creation of derivative products related to those publications. However, the raw imagery cannot be shared publicly. To adhere to these guidelines, we provide the data in HDF5 format, with the satellite imagery pre-processed to produce a derived product represented as a numpy array from Raster Vector. This process removes all geographic metadata.This data is for academic use only and should not be used commercially. Proper credit to the current authors, Planet Labs [16], and the Alberta Energy Regulator [30] is required when using this data.
Appendix D Dataset Samples Illustration

### v2

@@ B.5 PlanetScope Satellite Imagery
B.6 Meta Data Description
@@
B.7 Label Data Description
For our experiments, we create single-channel segmentation maps, which are binary maps used to locate instances of wells. We also generate multi-class segmentation maps, where each class denotes a well in an active, abandoned, or suspended state. Furthermore, we provide COCO format object detection labels for wells. In both segmentation and detection labels, we represent various states with class IDs as ’Active’: 1, ’Suspended’: 2, ’Abandoned’: 3. To maintain consistency, we standardize the diameter of a well site to 90 meters (typically ranging from 70 to 120 meters) when annotating, resulting in a 30-pixel diameter in the labels. Figure LABEL:fig:figure1a(b) illustrates image patches with their corresponding labels, Figure LABEL:fig:figure1a(a) illustrates various spectral bands present in an image alongside the original image with bounding box annotations for reference and an example of a bounding box label in COCO format is shown below.
@@ B.7 Label Data Description
Appendix C Dataset Samples Illustration
Samples from the dataset, covering various scenarios, are shown in Figures 13 and 14.